OpenStudy (anonymous):

Which of the following is a solution of x2 + 5x = -2?

OpenStudy (satan ):

Set it all equal to 0.

OpenStudy (anonymous):

ok i did that and i got \[5+ or - \sqrt{33}/ 2\]

OpenStudy (anonymous):

@Satan 

OpenStudy (satan ):

Using the quadratic formula right?

OpenStudy (anonymous):

yeah and i got that as my final answer. im not sure if thats correct

OpenStudy (satan ):

Well you're close.. you added 8 instead of subtracting. And it's \(-5 \pm\)

OpenStudy (satan ):

Because remember, when you get it in standard form it'll be \(0=x^2+5x+2\) And you plug those into the formula like so \(\Large \frac{-(5) \pm \sqrt{(5)^2-4(1)(2)}}{2(1)}\)

OpenStudy (anonymous):

ok then im pretty sure its this \[-5 + or - \sqrt{17}/2\]

OpenStudy (satan ):

Yup, there you go!

OpenStudy (anonymous):

wow thank you!

OpenStudy (satan ):

You're welcome :)